OpenStudy (anonymous):

Can anyone please help me with this question? A ball is dropped from the top of a 1,000 foot building. The height of the ball is half its original height after each bounce. What will the height of the ball be after 12 bounces? Using complete sentences, explain the procedure taken to answer this question.

OpenStudy (anonymous):

The height after the first bounce is 1000/2 = 500 ft. After the second bounce it is (1000/2)/2 = 250 ft. Third bounce, ((1000/2)/2)/2 = 125 ft. and so on. This can be generalized to Max height = 1000 / 2^n where n is the number of bounces. With n = 10, this becomes 1000 / 2^10 = 1000/1024 = 0.9765625 ft.